# A Survey on Medical Large Language Models: Technology, Application, Trustworthiness, and Future Directions

### v1

@@ IV-F2 Specific Multi-Modal (Medical) LLMs
AD-MM-LLM [62]. For diagnosing Alzheimer’s disease (AD), the study [62] introduced a pre-trained LLM to non-image data for knowledge embedding and a ConvNeXt for image data. A multi-modal alignment is conducted for multi-level multi-modal feature fusion.
RAMM [63]. RAMM is a retrieval-augmented pretrain-and-finetune paradigm [63], which could alleviate the data limitation issue for biomedical VQA. A new biomedical dataset named PMCPM is constructed by extracting image-text pairs from PubMed. Image-text contrastive objective (ITC) is exploited for pre-training. A retrieval-augmented method is to retrieve similar image-text pairs based on ITC from pre-training datasets.
LLaVA-Med [40]. LLaVA-Med exhibits excellent multi-modal conversational capability. LLaVA-Med [40] is required to align biomedical vocabulary based on figure-caption pairs and learn open-ended semantics using generated instruction-following data, which can follow open-ended instructions to assist with queries about a biomedical image. The training data includes a biomedical figure-caption dataset extracted from PubMed Central. Besides, LLaVA-Med can be trained in less than 15 hours using 8 A100s.
Qilin-Med-VL [64]. Qilin-Med-VL [64] is the first Chinese large vision-language model designed for text and visual data, which relies on a pre-trained vision transformer with a foundational LLM. The training dataset is ChiMed-VL with more than 1M image-text pairs. A two-stage curriculum training process is conducted including feature alignment and instruction tuning, which improves the ability to generate medical captions and answer complex medical queries.
@@ V Applying Medical LLMs
This section delves into a comprehensive investigation of the multifarious applications of LLMs in the medical domain. Med-LLMs are experiencing rapid growth in the clinical context, medical education, and medical research, including clinical decision, medical record analysis, patient engagement, health information dissemination, etc.
@@ V-A Multifarious Applications
V-A1 Clinical Decision Support
Med-LLMs can play a crucial role in enhancing clinical decision support systems [69], which are designed to assist healthcare professionals in making informed decisions about patient care. Here are some specific ways Med-LLMs contribute to clinical decision support: (1) Symptom Analysis [70]: By understanding natural language descriptions of symptoms entered by healthcare providers or patients, Med-LLMs can analyze and correlate these symptoms with a vast database of medical knowledge to suggest possible diagnoses. This can be particularly helpful in identifying complex or rare conditions. (2) Risk Assessment: Med-LLMs can evaluate patient data, including demographics, medical history, lifestyle factors, and current health status, to assess the risk of developing certain conditions or complications. This enables proactive interventions and personalized prevention strategies [223]. (3) Treatment Recommendations [224, 225, 226]: Based on the diagnosed condition, Med-LLMs can propose evidence-based treatment options, considering factors such as patient allergies, medication interactions, and the latest clinical guidelines. This supports clinicians in selecting the most appropriate therapy for each patient.
@@ V-A2 Clinical Report Generation
Med-LLMs are also being leveraged for the generation of clinical reports [229], streamlining documentation workflows and improving the accuracy and consistency of medical reporting. Med-LLMs can provide a flexible way for clinical report generation: (1) Automated Report Writing [110]: By processing input data from various sources such as electronic health records (EHRs), imaging study results, and laboratory tests, Med-LLMs can generate comprehensive, structured clinical reports automatically. This includes discharge summaries, radiology reports, pathology reports, and operative notes, saving time for healthcare professionals. (2) Customizable Templates [230]: Med-LLMs can incorporate institutional or specialty-specific templates, ensuring that generated reports follow the desired format and include all necessary information. This customization enhances the utility and readability of the reports for different clinical contexts. (3) Consistency and Accuracy: By relying on a standardized knowledge base, Med-LLMs can reduce variability in language and improve the overall quality of reports. They ensure adherence to clinical standards and minimize the risk of errors or inconsistencies that can arise from manual reporting.
V-A3 Medical Education Training
@@
Med-LLMs are transforming medical education by offering innovative ways to teach, learn, and interact with medical knowledge [231]. Several key directions of Med-LLMs in the field of medical education involve (1) Question-Answering Systems: Students and practitioners can pose complex medical questions to these models, receiving accurate and detailed explanations drawn from a wide range of medical literature. This fosters a deeper understanding of concepts and encourages critical thinking. (2) Simulation of Clinical Cases [32]: Med-LLMs can simulate virtual patients with specific symptoms, medical histories, and test results, enabling learners to practice diagnostic reasoning and treatment planning in a safe, controlled environment [232, 233]. (3) Summarization of Research Papers [234]: Given the vast amount of medical literature published daily, LLMs can summarize key findings, methodologies, and implications from research papers, helping students stay updated with the latest advancements efficiently.
V-A4 Medical Assistive Robotics
@@
The integration of Med-LLMs with Medical Robotics is an emerging field that promises to revolutionize surgical training, patient care, and the development of advanced robotic systems. Here are some key areas where Med-LLMs can contribute to medical robotics: (1) Patient-Specific Surgical Planning [235]: With access to a patient’s medical records and imaging data, Med-LLMs can assist in creating personalized surgical plans, taking into account the unique anatomy and health status of each patient. This information can guide robotic systems in executing highly customized surgeries. (2) Robotic Surgical Training [236]: By generating realistic surgical scenarios and providing real-time feedback, Med-LLMs can enhance the training of surgeons using robotic platforms. They can simulate various complications and patient responses, creating a more comprehensive and dynamic learning experience.
V-A5 Medical Language Translation
@@
Med-LLMs can overcome linguistic barriers to enhance communication and improve healthcare delivery across diverse populations [237]. Here’s how Med-LLMs are revolutionizing the medical language translation task: (1) Precision in Medical Terminology: Trained on extensive medical corpora, these models grasp the intricacies of medical vocabulary, ensuring translations are not just linguistically accurate but also contextually precise. This is crucial for avoiding misunderstandings that could lead to misdiagnosis or incorrect treatment. (2) Real-Time Translation [238]: In clinical settings, Med-LLMs can facilitate real-time conversations between healthcare providers and patients speaking different languages. This live translation capability streamlines communication during consultations, emergency situations, or remote care scenarios. (3) Patient Documentation [239]: They can automatically translate patient records, intake forms, discharge summaries, and consent forms, ensuring that medical staff can access and understand crucial information promptly, regardless of the original language
@@
It’s crucial to note that while these tools can greatly enhance the efficiency and accuracy of clinicians, they should not replace human expertise and clinical judgment. Med-LLMs are meant to augment rather than supplant the role of healthcare providers, ensuring that diagnoses are made with appropriate context and ethical considerations. Direct patient care and final diagnostic decisions should always be under the supervision of qualified medical professionals.
@@ V-B Unique Challenges for Med-LLMs
Protected Health Information. In the construction of digital medical infrastructure, implementing stringent security protocols and privacy measures is crucial. Healthcare entities bear significant responsibility to ensure that patients have confidence in the management, transmission, and use of their health data. Furthermore, these entities must adhere to relevant legal and ethical standards to strictly protect the confidentiality of patient information, thereby maintaining the credibility and reliability of medical services [240, 79].
Clinical Workflows. LLMs demonstrate significant potential in areas such as education, laboratory, and pathology analysis [241], and facilitating Shared Decision Making (SDM) [242]. However, integrating LLMs into existing hospital workflows presents challenges. This integration process requires meticulous planning to adapt to specific medical environments and to overcome technological, ethical, and operational barriers. Only through careful planning and effective management of these challenges can LLMs be seamlessly incorporated into hospital workflows, thereby enhancing the efficiency and quality of patient care [243].
Safety and Accountability. Due to their “black box” nature, LLMs present complex challenges in interpretability. Therefore, to ensure the accountability and safety of LLMs, stringent supervision and verification measures must be implemented. These measures are designed to maintain the integrity and credibility of LLM applications, thereby ensuring their effectiveness and reliability in practical use [79].
@@ VI Trustworthiness and Safety
VI-A Fairness
Fairness, a concept stemming from sociology, economics, and law, is described by the Oxford English Dictionary as “imperfect and just treatment or behavior without favoritism or discrimination” [65, 244]. In NLP, ensuring fairness requires the elimination of social biases embedded in the model, as these biases can lead to unfair treatment of specific groups [245] in the process of model coding and task processing. Therefore, a fair language model must be free from bias, which makes fairness and social bias intertwined in NLP research.
@@ VI-A Fairness
Research has revealed biases in healthcare [72], and LLMs are able to capture these biases from training data [65, 246, 247, 248] and amplify existing biases [249, 250], producing medically unfair outputs. Given the various biases (e.g., race, gender, and disability) present in clinical practice and research corpora used for training large language models [72, 251, 80], and the potential harms that unfair large language model systems can bring to medicine [252], ensuring the fairness of large language models in medical applications is crucial. The medical bias generated by large models can come from skewed data and model limitations. If biased training data is used, LLMs may retain or even enhance such bias [71, 79, 80]. At the same time, bias also comes from the limitations of models such as design specifications, structures, and algorithms [253].
At present, there are various technologies for quantifying LLM bias in the medical field. Among them, disparity metrics play an important role in bias detection. Taking BiasMedQA [254] as an example, it is a benchmark for evaluating cognitive bias in medical tasks, which can effectively quantify the specific levels of cognitive bias in six LLMs (GPT-4, Mixtral-8x70B, GPT-3.5, PaLM-2, Llama 2 70B-chat, and PMC Llama 13B focusing on the medical field). In addition, counterfactual evaluation can provide insight into the model’s understanding of causal relationships by constructing hypothetical scenarios, thereby revealing and quantifying potential biases in the medical field [255, 256].
@@
Current strategies to eliminate LLM bias in the medical field are also diversifying. For open-sourced LLMs, reinforcement learning using feedback from clinicians, etc. [257] is a promising way to counteract bias. For non-open source LLMs, it is often necessary to design debiasing programs according to the downstream task training strategy. The medium-sized LLMs previously developed by GPT-3 use pre-training and fine-tuning paradigms as the main training strategies and such LLMs’ social bias can be debiased from a variety of perspectives. For example, the data augmentation technique  [258] is used to optimize the training data during the pre-training phase to effectively reduce bias. LLMs with 100 million parameters are developing rapidly based on demonstrations. Since the representations of most closed-source large-size LLMs are not available, it is more difficult to eliminate bias in the responses of large-scale LLMs, and the commonly used methods include instruction fine-tuning  [38] and prompt engineering  [254, 259] strategies.
VI-B Accountability
Lack of accountability in LLMs is recognized as an obstacle hindering its application in the medical field [66]. To address this challenge, clinical practice, and research ensure the reliability of LLMs through interpretive methods and human oversight procedures.
@@ VI-B Accountability
Interpretive methods increase the transparency of the model and help people understand its decision-making process. Amir Hassan Shariatmadari et al.’s research [260] enhances the interpretability of LLMs in the biomedical field by integrating knowledge graphs with large language models and visualizing attention probabilities. In addition to attention visualization, some recent studies [261] have used Local explanation models to assist LLMs and improve the interpretability of LLMs in the medical field. Human oversight procedures are crucial for ensuring transparency and reliability in LLM operations, encompassing clinical trial protocol development [262] and continuous monitoring of model performance [71, 251], thereby involving human experts to bridge the accountability gap and improve LLMs’ effectiveness in medical scenarios.
VI-C Privacy
In the medical field, data privacy is an important challenge for LLMs-based applications. LLMs need to process a large and diverse dataset during training, which can lead to the model inadvertently capturing and copying sensitive information, which can lead to privacy leakage during text generation  [67]. In addition, issues such as unintentional data memory, data leakage, and potential leakage of personal information have become the main privacy challenges faced by LLMs [263].
@@ VI-C Privacy
At present, privacy protection for privacy leakage and passive privacy attacks can be divided into pre-training and fine-tuning protection methods according to the location of privacy protection [73]. De-identification technology protects privacy by performing thorough data processing of sensitive information during the pre-training phase. DeID-GPTcite [158] uses the powerful Named Entity Recognition (NER) capability of large models to identify sensitive information, to achieve automatic de-identification, showing high accuracy and significant reliability, and is one of the earliest studies using LLM for medical text data processing and de-identification. At the same time, MA Rahman [264] proposed that federated learning [265], differential privacy [266] can reduce data breaches in the healthcare industry due to LLM API calls during fine-tuning.
VI-D Robustness
One of the possible future research directions of Med-LLMs is to explore the construction of effective adversarial test samples in the medical field, including the construction of synthetic anomaly cases  [267] and boundary stress testing  [268] to evaluate the robustness of large language models in the medical field. Due to the limited research in this area, especially as Alberts pointed out in the 2023 study  [68], there are many challenges in constructing medically relevant adversarial test samples, so future research should focus on overcoming these challenges and developing adversarial test methods suitable for the medical field.
@@ VI-D Robustness
There have been many studies that use uncertainty quantification methods to improve robustness. For example, Ke Shen et al. [269] proposed a formalism and method to improve the robustness of large language models by using risk-adjusted confidence scores. Future research can continue to explore how to apply uncertainty quantification techniques to large language models in the medical field to improve the robustness of the model and reduce risks. At the same time, considering the difficulties of large language models in processing out-of-distribution inputs, future research can also focus on how to use techniques similar to LLM-TTA  [270] to improve the robustness of the model to unknown or rare cases in the medical field. By effectively using the augmentation data generated by the model as a means of augmentation at test time, we can reduce the dependence of the model on expensive augmentation while maintaining the performance of the model, which will have important implications for advancing the application of large language models in the medical field.

### v2

@@ IV-F2 Specific Multi-Modal (Medical) LLMs
$\bullet$ AD-MM-LLM [62]. For diagnosing Alzheimer’s disease (AD), the study [62] introduced a pre-trained LLM to non-image data for knowledge embedding and a ConvNeXt for image data. A multi-modal alignment is conducted for multi-level multi-modal feature fusion.
$\bullet$ RAMM [63]. RAMM is a retrieval-augmented pretrain-and-finetune paradigm [63], which could alleviate the data limitation issue for biomedical VQA. A new biomedical dataset named PMCPM is constructed by extracting image-text pairs from PubMed. Image-text contrastive objective (ITC) is exploited for pre-training. A retrieval-augmented method is to retrieve similar image-text pairs based on ITC from pre-training datasets.
$\bullet$ LLaVA-Med [40]. LLaVA-Med exhibits excellent multi-modal conversational capability. LLaVA-Med [40] is required to align biomedical vocabulary based on figure-caption pairs and learn open-ended semantics using generated instruction-following data, which can follow open-ended instructions to assist with queries about a biomedical image. The training data includes a biomedical figure-caption dataset extracted from PubMed Central. Besides, LLaVA-Med can be trained in less than 15 hours using 8 A100s.
$\bullet$ Qilin-Med-VL [64]. Qilin-Med-VL is the first Chinese large vision-language model designed for text and visual data, which relies on a pre-trained vision transformer with a foundational LLM. The training dataset is ChiMed-VL with more than 1M image-text pairs. A two-stage curriculum training process is conducted including feature alignment and instruction tuning, which improves the ability to generate medical captions and answer complex medical queries.
@@ V Applying Medical LLMs
This section delves into a comprehensive investigation of the multifarious applications of LLM in the medical domain. Med-LLMs are experiencing rapid growth in the clinical context, medical education, and medical research, including clinical decision, medical record analysis, patient engagement, health information dissemination, etc.
@@ V-A Multifarious Applications
$\bullet$ Clinical Decision Support. Med-LLMs can play a crucial role in enhancing clinical decision support systems [69], which are designed to help healthcare professionals make informed decisions about patient care. Here are some specific ways Med-LLMs contribute to clinical decision support: (1) Symptom Analysis [70]: By understanding natural language descriptions of symptoms entered by healthcare providers or patients, Med-LLMs can analyze and correlate these symptoms with a vast database of medical knowledge to suggest possible diagnoses. This can be particularly helpful in the identification of complex or rare conditions. (2) Risk Assessment: Med-LLMs can assess patient data, including demographics, medical history, lifestyle factors, and current health status, to assess the risk of developing certain conditions or complications. This enables proactive interventions and personalized prevention strategies [217]. (3) Treatment Recommendations [218, 219, 220]: Based on the diagnosed condition, Med-LLMs can propose evidence-based treatment options, considering factors such as patient allergies, medication interactions, and the latest clinical guidelines. This supports clinicians in selecting the most appropriate therapy for each patient.
@@
$\bullet$ Clinical Report Generation. Med-LLMs are also being leveraged for the generation of clinical reports [223], streamlining documentation workflows and improving the accuracy and consistency of medical reporting. Med-LLMs can provide a flexible way for clinical report generation: (1) Automated Report Writing [94]: By processing input data from various sources such as electronic health records (EHRs), imaging study results, and laboratory tests, Med-LLMs can generate comprehensive, structured clinical reports automatically. This includes discharge summaries, radiology reports, pathology reports, and operative notes, saving time for healthcare professionals. (2) Customizable Templates [224]: Med-LLMs can incorporate institutional or specialty-specific templates, ensuring that generated reports follow the desired format and include all necessary information. This customization enhances the utility and readability of the reports for different clinical contexts. (3) Consistency and Accuracy: By relying on a standardized knowledge base, Med-LLMs can reduce variability in language and improve the overall quality of reports. They ensure adherence to clinical standards and minimize the risk of errors or inconsistencies that can arise from manual reporting.
@@
$\bullet$ Medical Education Training. Med-LLMs are transforming medical education by offering innovative ways to teach, learn, and interact with medical knowledge [225]. Several key directions of Med-LLMs in the field of medical education involve (1) Question-Answering Systems: Students and practitioners can pose complex medical questions to these models, receiving accurate and detailed explanations drawn from a wide range of medical literature. This fosters a deeper understanding of concepts and encourages critical thinking. (2) Simulation of Clinical Cases [32]: Med-LLMs can simulate virtual patients with specific symptoms, medical histories, and test results, enabling learners to practice diagnostic reasoning and treatment planning in a safe, controlled environment [226, 227]. (3) Summarization of Research Papers [228]: Given the vast amount of medical literature published daily, LLMs can summarize key findings, methodologies, and implications from research papers, helping students stay updated with the latest advancements efficiently.
@@
$\bullet$ Medical Assistive Robotics. The integration of Med-LLMs with Medical Robotics is an emerging field that promises to revolutionize surgical training, patient care, and the development of advanced robotic systems. Here are some key areas where Med-LLMs can contribute to medical robotics: (1) Patient-Specific Surgical Planning [229]: With access to a patient’s medical records and imaging data, Med-LLMs can assist in creating personalized surgical plans, taking into account the unique anatomy and health status of each patient. This information can guide robotic systems in executing highly customized surgeries. (2) Robotic Surgical Training [230]: By generating realistic surgical scenarios and providing real-time feedback, Med-LLMs can enhance the training of surgeons using robotic platforms. They can simulate various complications and patient responses, creating a more comprehensive and dynamic learning experience.
@@
$\bullet$ Medical Language Translation. Med-LLMs can overcome linguistic barriers to enhance communication and improve healthcare delivery across diverse populations [231]. Here’s how Med-LLMs are revolutionizing the medical language translation task: (1) Precision in Medical Terminology: Trained on extensive medical corpora, these models grasp the intricacies of medical vocabulary, ensuring translations are not just linguistically accurate but also contextually precise. This is crucial for avoiding misunderstandings that could lead to misdiagnosis or incorrect treatment. (2) Real-Time Translation [232]: In clinical settings, Med-LLMs can facilitate real-time conversations between healthcare providers and patients speaking different languages. This live translation capability streamlines communication during consultations, emergency situations, or remote care scenarios. (3) Patient Documentation [233]: They can automatically translate patient records, intake forms, discharge summaries, and consent forms, ensuring that medical staff can access and understand crucial information promptly, regardless of the original language
$\bullet$ Drug Discovery. The integration of Large Language Models (LLMs) into drug discovery marks a significant advancement in pharmaceutical research. These AI-driven models, trained on extensive biomedical data, expedite literature review and data mining, uncovering novel drug targets and disease mechanisms. LLMs improve computational drug development by predicting molecular properties such as solubility and toxicity and designing optimized compounds for therapeutic efficacy. They play a crucial role in virtual screening, efficiently identifying promising drug candidates from vast chemical libraries. In addition, LLMs contribute to the design and analysis of clinical trials, optimizing patient recruitment and treatment personalization. As LLMs continue to evolve, they promise to accelerate the discovery of innovative therapies, heralding a new era of precision medicine.
@@
It’s crucial to note that while these tools can greatly enhance the efficiency and accuracy of clinicians, they should not replace human expertise and clinical judgment. Med-LLMs are meant to augment rather than replace the role of healthcare providers, ensuring that diagnoses are made with appropriate context and ethical considerations. Direct patient care and final diagnostic decisions should always be under the supervision of qualified medical professionals.
@@ V-B Unique Challenges for Med-LLMs
$\bullet$ Protected Health Information. In the construction of digital medical infrastructure, implementing stringent security protocols and privacy measures is crucial. Healthcare entities bear significant responsibility to ensure that patients have confidence in the management, transmission, and use of their health data. Furthermore, these entities must adhere to relevant legal and ethical standards to strictly protect the confidentiality of patient information, thereby maintaining the credibility and reliability of medical services [234, 79].
$\bullet$ Clinical Workflows. LLMs demonstrate significant potential in areas such as education, laboratory, and pathology analysis [235], and facilitating Shared Decision Making (SDM) [236]. However, integrating LLMs into existing hospital workflows presents challenges. This integration process requires meticulous planning to adapt to specific medical environments and to overcome technological, ethical, and operational barriers. Only through careful planning and effective management of these challenges can LLMs be seamlessly incorporated into hospital workflows, thereby enhancing the efficiency and quality of patient care [237].
$\bullet$ Safety and Accountability. Due to their “black box” nature, LLMs present complex challenges in interpretability. Therefore, to ensure the accountability and safety of LLMs, stringent supervision and verification measures must be implemented. These measures are designed to maintain the integrity and credibility of LLM applications, thereby ensuring their effectiveness and reliability in practical use [79].
@@ VI Trustworthiness and Safety
$\bullet$ Fairness. As a concept stemming from sociology, economics, and law, fairness is described by the Oxford English Dictionary as “imperfect and just treatment or behavior without favoritism or discrimination” [65, 238]. In NLP, ensuring fairness requires the elimination of social biases embedded in the model, as these biases can lead to unfair treatment of specific groups [239] in the process of model coding and task processing. Therefore, a fair language model must be free from bias, which makes fairness and social bias intertwined in NLP research. Research has revealed biases in healthcare [72], and LLMs are able to capture these biases from training data [65, 240, 241, 242] and amplify existing biases [243, 244], producing medically unfair outputs. Given the various biases (e.g., race, gender, and disability) present in clinical practice and research corpora used for training large language models [72, 245, 80], and the potential harms that unfair large language model systems can bring to medicine [246], ensuring the fairness of large language models in medical applications is crucial. The medical bias generated by large models can come from skewed data and model limitations. If biased training data is used, LLMs may retain or even enhance such bias [71, 79, 80]. At the same time, bias also comes from the limitations of models such as design specifications, structures, and algorithms [247].
@@
At present, there are various technologies for quantifying LLM bias in the medical field. Among them, disparity metrics play an important role in bias detection. Taking BiasMedQA [248] as an example, it is a benchmark for evaluating cognitive bias in medical tasks, which can effectively quantify the specific levels of cognitive bias in six LLMs (GPT-4, Mixtral-8x70B, GPT-3.5, PaLM-2, Llama 2 70B-chat, and PMC Llama 13B focusing on the medical field). In addition, counterfactual evaluation can provide insight into the model’s understanding of causal relationships by constructing hypothetical scenarios, thereby revealing and quantifying potential biases in the medical field [249, 250]. Current strategies to eliminate LLM bias in the medical field are also diversifying. For open-sourced LLMs, reinforcement learning using feedback from clinicians, etc. [251] is a promising way to counteract bias. For non-open source LLMs, it is often necessary to design debiasing programs according to the downstream task training strategy. The medium-sized LLMs previously developed by GPT-3 use pre-training and fine-tuning paradigms as the main training strategies and such LLMs’ social bias can be debiased from a variety of perspectives. For example, the data augmentation technique  [252] is used to optimize the training data during the pre-training phase to effectively reduce bias. LLMs with 100 million parameters are developing rapidly based on demonstrations. Since the representations of most closed-source large-size LLMs are not available, it is more difficult to eliminate bias in the responses of large-scale LLMs, and the commonly used methods include instruction fine-tuning  [38] and prompt engineering  [248, 253] strategies.
@@
$\bullet$ Accountability. Lack of accountability in LLMs is recognized as an obstacle hindering its application in the medical field [66]. To address this challenge, clinical practice, and research ensure the reliability of LLMs through interpretive methods and human oversight procedures. Interpretive methods increase the transparency of the model and help people understand its decision-making process. [254] enhanced the interpretability of LLMs in the biomedical field by integrating knowledge graphs with large language models and visualizing attention probabilities. In addition to attention visualization, some recent studies [255] have used Local explanation models to assist LLMs and improve the interpretability of LLMs in the medical field. Human oversight procedures are crucial for ensuring transparency and reliability in LLM operations, encompassing clinical trial protocol development [256] and continuous monitoring of model performance [71, 245], thereby involving human experts to bridge the accountability gap and improve LLMs’ effectiveness in medical scenarios.
@@
$\bullet$ Privacy. Data privacy is an important challenge for medical applications. LLMs inadvertently capture sensitive information, leading to privacy leakage during text generation  [67]. Unintentional data memory, data leakage, and potential leakage of personal information have become the main privacy challenges faced by LLMs [257]. At present, privacy protection for privacy leakage and passive privacy attacks can be divided into pre-training and fine-tuning protection methods according to the location of privacy protection [73]. De-identification technology protects privacy by performing thorough data processing of sensitive information during the pre-training phase. DeID-GPTcite [150] uses the powerful Named Entity Recognition (NER) capability of large models to identify sensitive information, to achieve automatic de-identification, showing high accuracy and significant reliability, and is one of the earliest studies using LLM for medical text data processing and de-identification. At the same time, MA Rahman [258] proposed that federated learning [259], differential privacy [260] can reduce data breaches in the healthcare industry due to LLM API calls during fine-tuning.
@@
$\bullet$ Robustness. One of the possible future research directions of Med-LLMs is to explore the construction of effective adversarial test samples in the medical field, including the construction of synthetic anomaly cases  [261] and boundary stress testing  [262] to evaluate the robustness of large language models in the medical field. Due to the limited research in this area, especially as Alberts pointed out in the 2023 study  [68], there are many challenges in constructing medically relevant adversarial test samples, so future research should focus on overcoming these challenges and developing adversarial test methods suitable for the medical field. There have been many studies that use uncertainty quantification methods to improve robustness. For example, Ke Shen et al. [263] proposed a formalism and method to improve the robustness of large language models by using risk-adjusted confidence scores. Future research can continue to explore how to apply uncertainty quantification techniques to large language models in the medical field to improve the robustness of the model and reduce risks. At the same time, considering the difficulties of large language models in processing out-of-distribution inputs, future research can also focus on how to use techniques similar to LLM-TTA  [264] to improve the robustness of the model to unknown or rare cases in the medical field. By effectively using the augmentation data generated by the model as a means of augmentation at test time, we can reduce the dependence of the model on expensive augmentation while maintaining the performance of the model, which will have important implications for advancing the application of large language models in the medical field.